# Pathwise uniform convergence of a full discretization for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise

### v4

@@ Abstract
This paper analyzes a full discretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. The discretization uses the Euler scheme for temporal discretization and the finite element method for spatial discretization. A key contribution of this work is the introduction of a novel stability estimate for a discrete stochastic convolution, which plays a crucial role in establishing pathwise uniform convergence estimates for fully discrete approximations of nonlinear stochastic parabolic equations. By using this stability estimate in conjunction with the discrete stochastic maximal $L^{p}$-regularity estimate, the study derives a pathwise uniform convergence rate that encompasses general spatial $L^{q}$-norms. Moreover, the theoretical convergence rate is verified by numerical experiments.
Keywords: stochastic Allen-Cahn equation, Euler scheme, the finite element method, discrete stochastic maximal $L^{p}$-regularity
@@ 1 Introduction
Assume that $(\Omega,\mathcal{F},\mathbb{P})$ is a given complete probability space with a filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ are defined. Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$. For each $t\in\mathbb{R}_{+}$, define $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ by
Let $\mathcal{O}$ be a bounded convex polyhedral domain in $\mathbb{R}^{3}$, and let $\Delta$ be the Laplace operator with homogeneous Dirichlet boundary condition. This paper considers the following stochastic Allen-Cahn equation:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,\\ y(0)=v,\end{cases}$ (1.1)
where $T\in(0,\infty)$ and the initial value $v$ are given, and the operator $F$ will be defined later.
@@
So far, the numerical analysis of the stochastic Allen-Cahn equation in the literature generally considers only the $L^{p}(\Omega;L^{2})$-norm error estimate for the fully discrete approximations. There is a lack of pathwise uniform convergence estimates and convergence estimates with the general spatial $L^{q}$-norms. It is worth noting that Bréhier et al. [4] derived some pathwise uniform convergence estimates for an explicit temporal splitting numerical scheme; however, their analysis may not be directly applicable to general convex bounded domains (see [4, pp. 2116]), and extending this analysis to the full discretization of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise is nontrivial.
@@
    $\mathbb{E}\max_{1\leqslant j\leqslant J}\Big{\lVert}{\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta_{h})^{k-j}P_{h}g(t)\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{q}}^{p}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}$
    for all $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$ with $p\in(2,\infty)$ and $q\in[2,\infty)$. This estimate is significant as it facilitates the demonstration of the pathwise uniform convergence of fully discrete approximations for nonlinear stochastic parabolic equations. While Van Neerven and Veraar [36] have established similar estimates for the Euler scheme, their approach is not directly applicable to deriving the above stability estimate, as detailed in Section 3.2.
  • For a fully discrete approximation of Eq. 1.1, using the Euler scheme for temporal discretization and the finite element method for spatial discretization, we establish the following error estimate:
@@
    for all $p\in(2,\infty)$, $q\in[2,\infty)$ and $\varepsilon>0$. Notably, even for the three-dimensional stochastic Allen-Cahn equation with additive noise, to our knowledge, analogous error estimates for fully discrete approximations utilizing the Euler scheme have yet to be discovered. This finding fills an important gap in the existing literature.
Our analysis builds upon the discrete stochastic maximal $L^{p}$-regularity estimate developed in [21, 23], which was inspired by the stochastic maximal $L^{p}$-regularity theory in [35]. A key role is also played by a stability estimate of a spatial semi-discretization from our prior work [22]. We impose relatively mild regularity conditions on the operator $F$ in Eq. 1.1, and although our analysis assumes that the initial value $v$ belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, it can be extended to cases where the initial value $v$ has lower regularity. Our work introduces a novel method for the numerical analysis of nonlinear stochastic parabolic equations, using the discrete stochastic maximal $L^{p}$-regularity estimate. This offers a fresh perspective compared to previous studies in the field.
@@
The remainder of this paper is organized as follows. In Section 2, we describe a full discretization of Eq. 1.1 and present a pathwise uniform convergence estimate. In Section 3, we establish a stability estimate for a discrete stochastic convolution, and prove the convergence estimate discussed earlier. Section 4 presents numerical experiments to validate the established pathwise uniform convergence. Finally, Section 5 provides a summary of the findings and contributions of this study.
@@ 2 Full discretization
We first introduce some notations. For any Banach space $X$, let $\gamma(H,X)$ denote the space of all $\gamma$-radonifying operators from $H$ to $X$; see [12, Chapter 9]. For any $p\in(1,\infty)$ and Banach space $X$, let $L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;X))$ and $L_{\mathbb{F}}^{p}(\Omega\times(0,T);X)$ denote the spaces of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega;L^{2}(0,T;X))$ and $L^{p}(\Omega\times(0,T);X)$, respectively. For ease of notation, we abbreviate the Lebesgue space $L^{q}(\mathcal{O})$ to $L^{q}$ for all $q\in[1,\infty]$. In addition, let $W_{0}^{1,\infty}(\mathcal{O})$ and $W^{2,\infty}(\mathcal{O})$ denote two standard Sobolev spaces; see, e.g., [33].
Next, we define the operator $F$ in Eq. 1.1 as follows. Let $(f_{n})_{n\in\mathbb{N}}$ denote a sequence of continuous functions from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$ such that for each $n\in\mathbb{N}$,
where $C_{F}$ is a positive constant independent of $n$, $x$, $y$, $x_{1}$, $x_{2}$, $y_{1}$ and $y_{2}$. With a little abuse of notation, for each $n\in\mathbb{N}$ we will use $f_{n}$ to denote a Nemytski operator defined by $f_{n}(v):=f_{n}(\cdot,v(\cdot))$, $\forall v\in L^{2}$. Let $\{\lambda_{n}\}_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying $\sum_{n\in\mathbb{N}}\lambda_{n}<\infty$. Finally, for any $v\in L^{2}$ define
$F(v):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(v),$
where $h_{n}\otimes f_{n}(v)\in\gamma(H,L^{2})$ is defined by
$\big{(}h_{n}\otimes f_{n}(v)\big{)}(w):=(w,h_{n})_{H}f_{n}(v),\quad\forall w\in H.$
By [12, Proposition 9.1.9] and (2.1), a straightforward calculation gives
$\lVert{F(v)}\rVert_{\gamma(H,L^{2})}\leqslant c\big{(}1+\lVert{v}\rVert_{L^{2}}\big{)},\quad\forall v\in L^{2},$
where $c$ is a positive constant depending only on $C_{F}$ in (2.1), $\sum_{n=0}^{\infty}\lambda_{n}$, and the volume of $\mathcal{O}$.
Now, we introduce the full discretization for Eq. 1.1. Let $J$ be a positive integer, and set $t_{j}:=j\tau$ for each $0\leqslant j\leqslant J$, where $\tau:=T/J$. Let $\mathcal{K}_{h}$ be a conforming and quasi-uniform triangulation of $\mathcal{O}$ consisting of three-dimensional simplexes, and let $h$ denote the maximum size of the elements in $\mathcal{K}_{h}$. Define
$X_{h}:=\left\{v_{h}\in C(\overline{\mathcal{O}})\;\middle|\;\text{ $v_{h}$ is linear on each $K\in\mathcal{K}_{h}$ and $v_{h}$ vanishes on $\partial\mathcal{O}$}\right\},$
where $C(\overline{\mathcal{O}})$ is the set of all continuous functions on $\overline{\mathcal{O}}$. Let $P_{h}$ be the $L^{2}$-orthogonal projection onto $X_{h}$, and for any $u_{h}\in X_{h}$ define $\Delta_{h}u_{h}\in X_{h}$ by
@@
where $\mu$ denotes the three-dimensional Lebesgue measure on $\mathcal{O}$. The full discretization of Eq. 1.1 is given as follows:
[…]
Finally, we establish the convergence of the discretization Eq. 2.2 in the following theorem; the proof is deferred to Section 3.
@@
Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ and $\tau\leqslant h^{2}$. Let $y$ and $(Y_{j})_{j=0}^{J}$ be the solutions of Eqs. 1.1 and 2.2, respectively. Then for all $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following error estimate holds:
@@
Remark 2.1.
Regrading the regularity of the solution $y$ of Eq. 1.1, we refer the reader to [24, 25, 26], where the authors have derived some regularity estimates with the spatial $L^{2}$-norm and the spatial $H^{1}$-norm. Furthermore, the authors in [22] have extended these results by establishing regularity estimates in terms of the general spatial $L^{q}$-norms; for example, under the condition $v\in L^{\infty}$ they have shown that
$y\in L^{p}(\Omega;C([0,T];L^{q})),\quad\forall p,q\in(2,\infty).$
@@
    The spatial accuracy of $O(h^{2-\epsilon})$ in Eq. 2.3 can be improved to $O(h^{2})$ by imposing stronger regularity conditions on the operator $F$. Specifically, if there exist an $\epsilon\in(0,1)$ and a positive constant $c$ such that
    $\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1+\epsilon,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{\dot{H}^{1+\epsilon,q}}\big{)}$
    for all $v\in\dot{H}^{1+\epsilon,q}$ with $q\in[2,\infty)$, then the improved accuracy can be achieved. The above space $\dot{H}^{1+\epsilon,q}$ is defined in the subsequent section.

### v5

@@ Abstract
This paper analyzes a full discretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. The discretization combines the Euler scheme for temporal approximation and the finite element method for spatial approximation. A pathwise uniform convergence rate is derived, encompassing general spatial $L^{q}$-norms, by using discrete versions of deterministic and stochastic maximal $L^{p}$-regularity estimates. Additionally, the theoretical convergence rate is validated through numerical experiments. The primary contribution of this work is the introduction of a technique to establish the pathwise uniform convergence of finite element-based full discretizations for nonlinear stochastic parabolic equations within the framework of general spatial $L^{q}$-norms.
Keywords: stochastic Allen-Cahn equation, Euler scheme, finite element method, discrete stochastic maximal $L^{p}$-regularity, pathwise uniform convergence
@@ 1 Introduction
Let $T\in(0,\infty)$ denote the terminal time, and let $\mathcal{O}$ be a bounded, convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$. Let $\Delta$ represent the Laplace operator subject to homogeneous Dirichlet boundary conditions. This paper studies the following stochastic Allen-Cahn equation:
$\begin{cases}\mathrm{d}y(t)=\left(\Delta y+y-y^{3}\right)(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),&t\in[0,T],\\ y(0)=v,\end{cases}$ (1.1)
where $v$ denotes the prescribed initial condition,a and $W_{H}$ represents an $H$-cylindrical Brownian motion. The diffusion coefficient $F$ will be precisely defined later.
@@
So far, the numerical analysis of the stochastic Allen-Cahn equation in the literature has generally considered only the $L^{p}(\Omega;L^{2})$-norm error estimates for fully discrete approximations. There is a lack of pathwise uniform convergence estimates and convergence estimates with general spatial $L^{q}$-norms. Notably, Bréhier et al. [5] derived some pathwise uniform convergence estimates for an explicit temporal splitting numerical scheme; however, their analysis might not be directly applicable to general domains (see [5, pp. 2116]). Extending this analysis to the full discretization of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise is nontrivial.
@@
    $\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\left\|\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta_{h})^{k-j}g_{h}(t)\,\mathrm{d}W_{H}(t)\right\|_{L^{q}}^{p}\right]\leqslant C\left\|g_{h}\right\|_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
    for all $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,X_{h}))$, where $p\in(2,\infty)$ and $q\in[2,\infty)$. Here, $X_{h}$ denotes the finite element space and $L^{q}$ denotes $L^{q}(\mathcal{O})$. This estimate is useful for demonstrating pathwise uniform convergence of fully discrete approximations based on finite elements for nonlinear stochastic parabolic equations, within the framework of general spatial $L^{q}$-norms. While van Neerven and Veraar [40] have developed analogous estimates for the Euler scheme, their methodology does not directly apply to deriving the aforementioned stability estimate, as detailed in Subsection 4.2.
  • For a fully discrete approximation of Eq. 1.1, using the Euler scheme for temporal discretization and the finite element method for spatial discretization, we establish the following pathwise uniform convergence estimate:
@@
    for all $p\in(2,\infty)$, $q\in[2,\infty)$ and $\varepsilon>0$. Notably, even for the three-dimensional stochastic Allen-Cahn equation with additive noise, to the best of our knowledge, such error estimates for fully discrete approximations using the Euler scheme have not been reported previously. This finding fills an important gap in the existing literature.
Our analysis is based on discrete deterministic and stochastic maximal $L^{p}$-regularity estimates. We impose relatively mild regularity conditions on the diffusion coefficient $F$ in Eq. 1.1. Although our analysis assumes that the initial value $v$ is deterministic and belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, it can be extended to cases where the initial value $v$ has lower regularity. In addition, the pathwise uniform convergence of numerical approximations for SPDEs within the framework of Banach spaces has recently attracted considerable attention; see, for example, [7, 8, 21, 40]. However, existing research has primarily focused on temporal semidiscretizations. This work introduces a new approach for establishing pathwise uniform convergence for finite element-based full discretizations of nonlinear stochastic parabolic equations, within the general spatial $L^{q}$-norms framework.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce the notational conventions used throughout this work and present a regularity result for the model problem. In Section 3, we describe a full discretization and provide a pathwise uniform convergence estimate. In Section 4, we establish a stability estimate for a discrete stochastic convolution and provide a proof for the previously mentioned convergence estimate. Section 5 presents numerical experiments to validate the theoretical result. Finally, Section 6 summarizes the findings and contributions of this study.
2 Mathematical Setup
For any two Banach spaces $E_{1}$ and $E_{2}$, we denote by $\mathcal{L}(E_{1},E_{2})$ the space of all bounded linear operators from $E_{1}$ to $E_{2}$, and abbreviate $\mathcal{L}(E_{1},E_{1})$ to $\mathcal{L}(E_{1})$. The symbol $I$ represents the identity map. Let $\mathcal{O}$ be a bounded, convex domain in $\mathbb{R}^{3}$ of class $\mathcal{C}^{2}$ (cf. Section 11.1 in Chapter 1 of [44]). For any $q\in[1,\infty]$, let $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ denote the usual Sobolev spaces (see, e.g., [39]), and we shall abbreviate $L^{q}(\mathcal{O})$ to $L^{q}$ for simplicity. Let $\Delta$ denote the Laplace operator subject to homogeneous Dirichlet boundary conditions. For $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the norm
$\lVert{u}\rVert_{\dot{H}^{\alpha,q}}:=\lVert{(-\Delta)^{\alpha/2}u}\rVert_{L^{q}},\quad\forall u\in\dot{H}^{\alpha,q}.$
We use $\dot{H}^{-\alpha,q^{\prime}}$ to denote the dual space of $\dot{H}^{\alpha,q}$, where $q^{\prime}$ is the conjugate exponent of $q$, i.e., $11/q+1/q^{\prime}=$ 1. For any $q\in(1,\infty)$, it is well-known (see, e.g., Theorem 16.14 of [44]) that $\dot{H}^{2,q}$ coincides with the intersection $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$, and is equipped with a norm equivalent to that of $W^{2,q}(\mathcal{O})$. Moreover, $\dot{H}^{1,q}$ is identical to $W_{0}^{1,q}(\mathcal{O})$, with equivalent norms.
Assume that $(\Omega,\mathcal{F},\mathbb{P})$ is a given complete probability space with a right-continuous filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ are defined. For any $0\leqslant s<t<\infty$ and for any $n\in\mathbb{N}$, the increment $\beta_{n}(t)-\beta_{n}(s)$ is independent of $\mathcal{F}_{s}$. Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$. Let the $H$-cylindrical Brownian motion be defined such that for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is given by
We use the symbol $\mathbb{E}$ to represent the expectation operator associated with the probability space $(\Omega,\mathcal{F},\mathbb{P})$. For any UMD Banach space $E$, let $\gamma(H,E)$ denote the space of all $\gamma$-radonifying operators from $H$ to $E$. For any $p\in[2,\infty)$, we denote by $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,E))$ and $L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,E)))$ the spaces of all $\mathbb{F}$-adapted $\gamma(H,E)$-valued processes in $L^{p}(\Omega\times(0,T);\gamma(H,E))$ and $L^{p}(\Omega;L^{2}(0,T;\gamma(H,E)))$, respectively. When $E$ is a Hilbert space, $\gamma(H,E)$ coincides with the standard Hilbert-Schmidt space from $H$ to $E$, and we have the celebrated Itô’s isometry:
$\mathbb{E}\left\lVert\int_{0}^{T}g(t)\,\mathrm{d}W_{H}(t)\right\rVert_{E}^{2}=\mathbb{E}\int_{0}^{T}\lVert{g(t)}\rVert_{\gamma(H,E)}^{2}\,\mathrm{d}t,$
for all $g\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,E))$. For an in-depth discussion of $\gamma$-radonifying operators, the reader is referred to Chapter 9 of [16]. An extensive treatment of stochastic integrals with respect to $W_{H}$ in UMD Banach spaces can be found in [41].
Next, we define the diffusion coefficient $F$ in Eq. 1.1 as follows. Let $(f_{n})_{n\in\mathbb{N}}$ denote a sequence of continuous functions from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$ such that for each $n\in\mathbb{N}$,
where $C_{F}$ is a positive constant independent of $n$, $x$, $y$, $x_{1}$, $x_{2}$, $y_{1}$ and $y_{2}$. Let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying the condition $\sum_{n\in\mathbb{N}}\lambda_{n}<\infty$. Finally, for any $u\in L^{2}$ define
$F(u):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(\cdot,u(\cdot)),$
where $h_{n}\otimes f_{n}(\cdot,u(\cdot))\in\gamma(H,L^{2})$ is defined by
$\big{(}h_{n}\otimes f_{n}(\cdot,u(\cdot))\big{)}(w):=(w,h_{n})_{H}f_{n}(\cdot,u(\cdot)),\quad\forall w\in H.$
By [16, Proposition 9.1.9] and (2.1), a straightforward calculation gives the following growth property:
$\lVert{F(u)}\rVert_{\gamma(H,L^{2})}\leqslant c\big{(}1+\lVert{u}\rVert_{L^{2}}\big{)},\quad\forall u\in L^{2},$
where $c$ is a positive constant independent of $u$.
We call a process $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ a mild solution to the model problem Eq. 1.1 if it satisfies almost surely the following equality for all $t\in[0,T]$:
$y(t)=e^{t\Delta}v+\int_{0}^{t}e^{(t-s)\Delta}\big{[}y(s)-y^{3}(s)\big{]}\,\mathrm{d}s+\int_{0}^{t}e^{(t-s)\Delta}F(y(s))\,\mathrm{d}W_{H}(s),$
where $e^{t\Delta}$, $t\in[0,T]$, denotes the analytic semigroup generated by the Laplace operator $\Delta$ with homogeneous Dirichlet boundary conditions, and $v$ is the given initial value. While the condition $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ can be relaxed (see, e.g., Chapter 6 of [34]), it suffices for our purposes. Following the proof of Theorem 3.1 in [28], we establish the following regularity result. For related regularity results, we also refer the reader to [32, Section 3] and [30].
Proposition 2.1.
Suppose the initial value $v$ belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Then, the model problem (1.1) admits a unique mild solution. Moreover, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the solution $y$ satisfies
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q})\cap L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})).$
@@ 3 Full Discretization
Let $J$ be a positive integer, and set $t_{j}:=j\tau$ for each $0\leqslant j\leqslant J$, where $\tau:=T/J$ is the time step. Let $\mathcal{K}_{h}$ be a conforming and quasi-uniform triangulation of $\mathcal{O}$ consisting of three-dimensional simplexes, and let the spatial mesh size $h$ denote the maximum diameter of the simplexes in $\mathcal{K}_{h}$. Define
$X_{h}:=\Big{\{}u_{h}\in C(\overline{\mathcal{O}})\,\big{|}\,u_{h}=0\text{ on }\partial\mathcal{O}_{h}\text{ and outside of }\mathcal{O}_{h},\text{ and }u_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\Big{\}},$
where $C(\overline{\mathcal{O}})$ is the set of all continuous functions on the closure of $\mathcal{O}$, $\mathcal{O}_{h}$ is the closure of the union of the simplexes in $\mathcal{K}_{h}$, $r\geqslant 1$ is an integer, and $P_{r}(K)$ is the set of all polynomials on $K$ with degree not exceeding $r$ for each $K\in\mathcal{K}_{h}$. Let $P_{h}$ be the $L^{2}$-orthogonal projection operator onto $X_{h}$, and, for any $u_{h}\in X_{h}$, define $\Delta_{h}u_{h}\in X_{h}$ by
@@
where $\mu$ denotes the three-dimensional Lebesgue measure on $\mathcal{O}$.
This study considers the following full discretization of the model problem Eq. 1.1:
$\begin{cases}Y_{j+1}-Y_{j}=\tau\left(\Delta_{h}Y_{j+1}+Y_{j}-P_{h}Y_{j+1}^{3}\right)+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})\,\mathrm{d}W_{H}(t),\quad 0\leqslant j<J,\\ Y_{0}=P_{h}v.\end{cases}$ (3.1)
For this full discretization, we establish the following pathwise uniform convergence estimate; the proof is deferred to Section 4.
@@
Assume that the initial value $v$ belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, and $\tau\leqslant h^{2}$. Let $y$ be the mild solution of the model problem Eq. 1.1, and let $(Y_{j})_{j=0}^{J}$ be the solution of the full discretization Eq. 3.1. Then for all $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following pathwise uniform convergence estimate holds:
@@
    Assume that the sequence $(f_{n})_{n\in\mathbb{N}}$, in addition to satisfying (2.1), consists of twice continuously differentiable functions and satisfies the following growth condition for each $n\in\mathbb{N}$:
    $\left|\nabla_{x}^{2}f_{n}(x,y)\right|+\left|\nabla_{x}\nabla_{y}f_{n}(x,y)\right|+\left|\nabla_{y}^{2}f_{n}(x,y)\right|\leqslant C_{F}(1+|y|),\quad\forall x\in\mathcal{O},\,\forall y\in\mathbb{R}^{3},$
    where $C_{F}$ is a positive constant independent of $n$. Under the condition that the initial value $v$ belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, the spatial accuracy $O(h^{2-\epsilon})$ in Eq. 3.2 can be improved to $O(h^{2})$ for all $p\in(2,\infty)$ and $q\in[2,\infty)$.